# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

### v2

@@ Abstract
Drones have become prevalent robotic platforms with significant potential in Embodied AI. A crucial capability for drone-based Embodied AI is Referring Expression Comprehension (REC), which enables locating objects with language expressions. Despite advances in REC for ground-level scenes, drones’ unique capability for broad observation introduces distinct challenges: multiple potential targets, small-scale objects, and complex environmental contexts. To address these challenges, we introduce RefDrone, an REC benchmark for drone scenes. RefDrone reveals three key challenges: 1) multi-target and no-target scenarios; 2) multi-scale and small-scale target detection; 3) complex environment with rich contextual reasoning. To efficiently construct this dataset, we develop RDAnnotator, a semi-automated annotation framework where specialized modules and human annotator collaborate through feedback loops. RDAnnotator ensures high-quality contextual expressions while reducing annotation costs. Furthermore, we propose Number GroundingDINO (NGDINO), a novel method to handle multi-target and no-target cases. NGDINO explicitly estimates the number of objects referred to in the expression and incorporates this numerical pattern into the detection process. Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on RefDrone, as well as on the general-domain gRefCOCO and remote sensing RSVG benchmarks. The data and code are completely available at https://github.com/sunzc-sunny/refdrone.
@@ 1 Introduction
Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue [1, 42, 45]. Their ability to move freely and to observe broadly make them important platforms for Embodied AI applications [12, 35, 23, 14, 31]. A crucial capability in Embodied AI is Referring Expression Comprehension (REC) [43, 48, 4, 49], which serves as a critical bridge between natural language understanding and visual perception. REC requires drones to localize specific objects in images based on natural language expressions. However, existing REC datasets primarily focus on ground-level perspectives, such as RefCOCO [54, 36] dataset. The application of REC in drone-based scenarios presents unique challenges.
In this work, we introduce RefDrone, a challenging REC benchmark for drone scenes. RefDrone comprises 17,900 referring expressions annotated over 8,536 images, with 63,679 object instances. As illustrated in Figure 1, RefDrone poses three primary challenges that distinguish it from existing datasets: (1) multi-target and no-target samples, where expressions may refer to any number of objects, ranging from 0 to 242; (2) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; and (3) complex environment with rich contextual reasoning, encompassing diverse viewpoints, lighting conditions, and intricate backgrounds. The referring expressions richly capture spatial relations, object attributes, and inter-object interactions. As summarized in Table 1, RefDrone offers greater diversity and complexity than previous REC benchmarks. We further evaluate 13 representative REC models using both fine-tuning and zero-shot settings. Notably, all models exhibit significantly poorer performance on RefDrone compared to standard REC datasets (e.g., GroundingDINO [34] yields 21.07 $Acc_{img.}$ on RefDrone vs. 91.4 $Acc_{img.}$ on RefCOCO), highlighting the inherent difficulty and unique challenges posed by our benchmark.
@@
To efficiently construct RefDrone, we develop RDAnnotator, a semi-automated annotation pipeline for referring expression annotation in drone scenes. RDAnnotator leverages multiple specialized LMM-based modules, which collaborate within a feedback loop to generate and validate annotations. This tool reduces human involvement to quality control and minor adjustments, thereby significantly lowering annotation costs, while maintaining high-quality standards for complex expressions. RDAnnotator achieves a cost of just \$0.0539 per expression (GPT-4o API usage) and reduces human annotation time to under one minute per expression. This cost-efficiency makes RDAnnotator a scalable solution for large-scale dataset construction and can be readily adapted to other REC tasks.
@@
Furthermore, we propose Number GroundingDINO (NGDINO) to address the challenges of multi-target and no-target referring expressions. Our key insight is that explicitly utilizing the number information of referred objects enhances the handling of these scenarios. NGDINO includes three key components: (1) a number prediction head that estimates the count of target objects, (2) learnable number-queries that capture numerical patterns corresponding to varying object quantities, and (3) a number cross-attention module that fuses number queries with detection queries to inform object localization. Extensive experiments on both our RefDrone benchmark, as well as the general-domain gRefCOCO [28] and remote sensing RSVG [56] datasets, demonstrate that NGDINO substantially improves performance, particularly in the challenging multi-target and no-target cases.
@@
  • RefDrone Benchmark: We introduce RefDrone, a comprehensive benchmark for referring expression comprehension in drone scenes. RefDrone poses three key challenges: multi-target/no-target scenarios, multi-scale/small-object detection, and complex contextual reasoning. We provide thorough baseline evaluations with 13 representative REC models.
  • RDAnnotator Framework: We propose RDAnnotator, a semi-automated annotation framework that significantly reduces human effort while ensuring high-quality referring expressions. This framework is scalable for large-scale dataset construction and generalizable beyond drone imagery.
  • NGDINO Method: We develop NGDINO, a novel approach that explicitly models the number of referred objects to address multi-target and no-target cases. NGDINO achieves state-of-the-art performance on RefDrone and shows consistent improvements on gRefCOCO and RSVG benchmarks.
Table 1: Comparison of REC datasets relevant to RefDrone. Avg. objects: mean objects per expression. Avg. length: mean words per expression. No target: expressions without referred objects. Expression type: method for expression generation. Small target: percentage of small-scale objects.
RefCOCO/+/g [54, 36] gRefCOCO [28] D3 [51] RIS-CQ [19] RSVG [56] RefDrone (Ours) Image source COCO [26] COCO [26] COCO [26] VG [22]+COCO [26] DIOR [24] VisDrone [60] Avg. objects 1.0 1.4 1.3 3.6 2.2 3.8 Avg. length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target 0/0/0% 0.1% 6.3% - 17.2% 31.1%
@@ 2.1 Referring expression understanding datasets
Referring expression understanding identifies regions in visual content from natural language expressions. The main subtasks are Referring Expression Comprehension (REC), which outputs detection bounding boxes, and Referring Expression Segmentation (RES), which outputs segmentation masks. Datasets in this domain show increasing visual and linguistic complexity. Early datasets such as ReferIt [21] and RefCOCO [54] are pioneering but limited to simple expressions and single-target scenarios. Later datasets [32, 9, 28, 46, 19, 51] introduce more challenges. gRefCOCO [28] supports multi-target expressions. OV-VG [46] enables open-vocabulary queries. D3 [51] and RIS-CQ [19] offer more complex expressions. Domain-specific datasets [10, 8, 56, 37, 39] address specialized applications: RSVG [56] for remote sensing, RAVAR [37] for video action recognition, RIO [39] for affordance detection. However, drone scenes remain unexplored in REC, motivating our RefDrone benchmark. Table 1 compares RefDrone with other related REC benchmarks.
Recent dataset creation approaches leverage large language models (LLMs) to reduce annotation costs. Works such as LLaVA [30] and Ferret [53] demonstrate the effectiveness of LLMs in generating instruction tuning data. Similarly, RIS-CQ [19] and RIO [39] employ LLMs to generate referring expressions. However, these approaches treat LLMs as text generators and may lack proper visual grounding. In contrast, we employ a semi-automated annotation framework with iterative feedback mechanisms throughout the annotation pipeline to reduce annotation costs and improve quality.
@@ 2.2 Referring expression comprehension methods
REC methods can be broadly categorized into large multimodal models (LMMs) and specialist models. Large Multimodal Models [30, 53, 5, 7, 3, 38, 47, 27, 57, 50] have recently been applied to REC tasks as part of evaluating their broader visual-language understanding capabilities. These models leverage extensive referring instruction tuning data to achieve competitive performance without task-specific architectural designs. Despite their broad capabilities, LMMs consistently struggle with small object detection, primarily due to their inherent input resolution constraints. This constraint forces image downsampling, resulting in the loss of fine-grained visual details.
Specialist models include two-stage and one-stage methods. Two-stage methods [18, 29, 16, 44, 15] typically approach REC as a ranking task: first generating region proposals through a detector, then ranking them based on their alignment with the language input. However, they are often criticized for slow inference speeds. In contrast, one-stage methods [20, 25, 33, 13, 52, 34] directly predict target regions guided by language input. These approaches leverage transformer to enable cross-modal interactions between visual and textual features. Among these methods, GroundingDINO (GDINO) [34] has gained widespread attention for its impressive results in REC tasks. Our work extends GDINO by incorporating explicit number modeling to handle multi-target and no-target scenarios, which are critical challenges in drone-based REC tasks.
@@ 3 RefDrone benchmark
Figure 2: The overview of the RefDrone annotation process with RDAnnotator. Multiple specialized LMM-based modules collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
@@ 3.1 Data source
The RefDrone benchmark is built upon VisDrone2019-DET [60], a large-scale high-quality drone-captured object detection dataset. Source images are collected across multiple scenarios, illumination conditions, and flying altitudes. To ensure meaningful visual content, we filter out images with fewer than three objects and objects with bounding box areas below 64 pixels. VisDrone2019-DET provides object categories and bounding box coordinates, which we convert to normalized center points (range 0-1). This approach reduces the input context for LMMs while preserving spatial information.
@@ 3.2 RDAnnotator for semi-automated annotation
RDAnnotator (referring drone annotator) is a semi-automated annotation framework that minimizes human effort while ensuring high quality. It integrates specialized modules powered by LMMs with human validation. As shown in Figure 2, RDAnnotator employs five structured steps.
Step 1: scene understanding. A GPT-4o-based scene parsing module generates three diverse textual descriptions per image. These captions establish spatial layouts, object relationships, and contextual foundations to guide subsequent referring expression generation.
Step 2: color categorization. Color attributes are crucial discriminative features in referring expressions. We implement a hybrid color extraction pipeline, which combines low-level CNN-based classification (WideResNet-101 classifier [55]) with high-level LMM semantic reasoning, ensuring robust color attribution in challenging aerial conditions.
Step 3: expression generation. To address the complexity of referring expression generation, we reformulate the referring expression generation task as an object grouping problem. The target of the module is to group semantically related objects and provide appropriate reasons for each group. The reasons serve as the referring expressions. Furthermore, a dynamic feedback loop triggers Step 2 (color categorization) when unseen colors are detected, ensuring color attribute consistency.
Step 4: quality evaluation. A validation module assesses the semantic accuracy and uniqueness of each object-expression pair, classifying each annotation as either correct (accurately and uniquely identifying the described objects) or incorrect (inaccurate or ambiguous, with an explanation of the identified issue). Annotations deemed correct proceed to Step 5 (human verification), while incorrect annotations trigger targeted feedback: those with semantic errors are returned to Step 3 (expression generation), and color-related inaccuracies are redirected to Step 2 (color categorization).
Figure 3: Object number distribution per expression in gRefCOCO [28] and RefDrone datasets.
Figure 4: Object size distribution histograms in RefDrone and gRefCOCO [28] datasets.
Step 5: human verification. Human annotators review annotation outputs in three tiers:
@@
  • Refinement required. Minor errors are corrected through human editing.
  • Significant issues. Poorly grounded or inconsistent annotations trigger full regeneration.
Annotations with significant issues re-enter the pipeline at Step 4 (quality evaluation). If an annotation repeatedly fails, it is classified as a no-target sample. These feedback loops ensure all annotations meet quality standards and that no-target expressions are contextually relevant to the image.
Each step utilizes LMMs through in-context learning with carefully designed task-specific prompts and examples (see Appendix A.8). The final annotation quality demonstrates 42% directly accepted samples, 47% requiring minor refinement, and 11% demanding full re-annotation. RDAnnotator reduces human annotation effort by 85%, decreasing annotation time from 7 to 1 minute per expression, while incurring only \$0.0539 per expression in GPT-4o API costs. This cost-performance balance makes the framework scalable for large-scale dataset creation while maintaining annotation quality.
@@ 3.3 Dataset analysis
The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances in 10 categories. We retain the original train, validation, and test splits from VisDrone2019-DET [60]. The average expression length is 9.0 words, and each expression refers to an average of 3.8 objects. Figure 1 highlights the three key challenges in RefDrone:
1) Multi-target and no-target samples. Unlike conventional REC datasets [54, 36] that primarily focus on single-object references, RefDrone includes 11,362 multi-target and 847 no-target expressions. The number of referred targets per expression ranges from 0 to 242. Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where the majority of expressions contain references to only one or two objects.
2) Multi-scale and small-scale target detection. Figure 4 presents the challenges of scale distribution: small-scale objects ($<32^{2}$ pixels) make up 31% of instances, normal-scale ($32^{2}$-$96^{2}$ pixels) 55%, and large-scale ($>96^{2}$ pixels) 14%. Compared to gRefCOCO [28], RefDrone has a notably higher variance in object scales, emphasizing the dataset’s multi-scale challenges.
3) Complex environment with rich contextual reasoning. The images present inherent complexity, including diverse viewpoints, varying lighting conditions, and complex background environments. The referring expressions go beyond simple object attributes (e.g., color, category) and spatial relationships (e.g., ‘left’, ‘near’), incorporating rich object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’). Figure 5 visualizes this linguistic richness through word clouds of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.
3.4 Dataset comparison
Table 1 compares RefDrone with existing REC datasets. RefDrone stands out for its higher average number of referred objects and use of LMM for expression generation. Its expressions offer richer contextual details than template-based or human-annotated expressions [59]. While RIS-CQ [19] uses LLM for expression generation, it lacks visual content, often producing linguistically complex but visually disconnected expressions. Similarly, RSVG [56] focuses on small targets but lacks expression quality. In contrast, RefDrone comprehensively presents three challenges, establishing it as a challenging benchmark in REC tasks.

### v3

@@ Abstract
Drones have become prevalent robotic platforms with significant potential in Embodied AI. A crucial capability for drone-based Embodied AI is Referring Expression Comprehension (REC), which enables locating objects with language expressions. Despite advances in REC for ground-level scenes, drones’ unique capability for broad observation introduces distinct challenges: multiple potential targets, small-scale objects, and complex environmental contexts. To address these challenges, we introduce RefDrone, an REC benchmark for drone scenes. RefDrone reveals three key challenges: 1) multi-target and no-target scenarios; 2) multi-scale and small-scale target detection; 3) complex environments with rich contextual reasoning. To efficiently construct this dataset, we develop RDAnnotator, a semi-automated annotation framework where specialized modules and human annotators collaborate through feedback loops. RDAnnotator ensures high-quality contextual expressions while reducing annotation costs. Furthermore, we propose Number GroundingDINO (NGDINO), a novel method to handle multi-target and no-target cases. NGDINO explicitly estimates the number of objects referred to in the expression and incorporates this numerical pattern into the detection process. Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on RefDrone, as well as on the general-domain gRefCOCO and remote sensing RSVG benchmarks.
@@ 1 Introduction
Drones/UAVs have become increasingly prevalent in both personal and professional applications, including entertainment, package delivery, traffic surveillance, and emergency rescue [adao2017hyperspectral, san2018uav, varga2022seadronessee]. Their mobility and broad observational capabilities make them promising platforms for Embodied AI applications [fan-etal-2023-aerial, liu2023aerialvln, lee2024citynav, gao2024embodied, liu2024coherent]. A fundamental capability in Embodied AI is Referring Expression Comprehension (REC) [sima2023embodied, Wang_2024_CVPR, cai2024bridging, wang2024polaris], which bridges natural language understanding and visual perception by localizing specific objects in images based on textual descriptions. However, existing REC datasets predominantly adopt ground-level perspectives (e.g., RefCOCO [refcoco]), leaving drone-based scenarios largely underexplored despite their unique challenges.
Table 1: Comparison of REC datasets relevant to RefDrone. Avg. objects: objects per expression. Avg. length: words per expression. No target: expressions without referred objects. Expression type: expression generation method. Small target: percentage of small-scale objects.
RefCOCO/+/g [refcoco, refcocog] gRefCOCO [grefcoco] D3 [xie2024described] RIS-CQ [riscq] RSVG [zhan2023rsvg] RefDrone (Ours) Image source COCO [coco] COCO COCO VG [vg]+COCO DIOR [dior] VisDrone [visdrone] Avg. objects 1.0 1.4 1.3 3.6 2.2 3.8 Avg. length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM Templated LVLM Small target 0/0/0% 0.1% 6.3% - 17.2% 31.1%
[…]
@@
To enable efficient dataset construction, we develop RDAnnotator, a semi-automated annotation pipeline for referring expression annotation in drone scenes. RDAnnotator leverages multiple specialized LVLM-based modules, which collaborate within a feedback loop to generate and validate annotations. By reducing human involvement to quality control and minor refinements, RDAnnotator achieves a cost of merely \$0.0539 per expression (GPT-4o API) and reduces human annotation time to under one minute per expression. This cost-efficiency makes RDAnnotator a scalable solution for large-scale dataset construction and can be readily adapted to other REC tasks.
@@
Furthermore, we propose Number GroundingDINO (NGDINO) to address multi-target and no-target challenges. Our key insight is that explicitly modeling the number of referred objects significantly enhances handling of these scenarios. NGDINO includes three components: (1) a number prediction head estimating target object counts, (2) learnable number-queries capturing numerical patterns for varying object quantities, and (3) a number cross-attention module fusing number queries with detection queries for enhanced localization. Extensive experiments on our RefDrone benchmark, as well as the general-domain gRefCOCO [grefcoco] and remote sensing RSVG [zhan2023rsvg] datasets, demonstrate that NGDINO achieves substantial improvements, particularly in multi-target and no-target cases.
@@
  • RefDrone Benchmark: We introduce RefDrone, a comprehensive REC benchmark for drone scenes featuring three key challenges: multi-target/no-target scenarios, multi-scale/small-object detection, and complex contextual reasoning. We provide thorough evaluations of 26 representative REC models.
  • RDAnnotator Framework: We propose RDAnnotator, a cost-effective semi-automated annotation framework that significantly reduces human effort while ensuring high-quality annotations. This framework is scalable for large-scale dataset construction and generalizable beyond drone imagery.
  • NGDINO Method: We develop NGDINO, a novel approach explicitly modeling object counts to address multi-target and no-target cases. NGDINO achieves state-of-the-art performance on RefDrone with consistent improvements on gRefCOCO and RSVG.
@@ 2.1 Referring expression understanding datasets
Referring expression understanding identifies visual regions corresponding to natural language expressions. Subtasks include Referring Expression Comprehension (REC), which predicts bounding boxes, and Referring Expression Segmentation (RES), which generates pixel-level masks. Early benchmarks such as ReferIt [kazemzadeh2014referitgame] and RefCOCO [refcoco] were pioneering but mostly limited to single-target scenarios. Subsequent works introduced greater complexity. For instance, gRefCOCO [grefcoco] expanded to multi-target expressions, while D3  [xie2024described] and RIS-CQ [riscq] provided more descriptive and challenging language. However, these datasets typically involve a small, fixed number of targets per expression.
Domain-specific extensions have also emerged, including video action recognition (RAVAR [peng2024ravar]) and affordance detection (RIO [qu2024rio]). In the aerial domain, RSVG [zhan2023rsvg] focuses on remote sensing data but uses expressions with limited contextual richness. AerialVG [Liu_2025_ICCV] focuses on using spatial relations for a single target, leaving the challenge of multi-target and no-target tasks. To fill this gap, our RefDrone benchmark provides a more complex and realistic setting by introducing complex multi-target scenarios with contextually rich expressions. Table 1 provides a detailed comparison with related benchmarks.
Figure 2: The overview of the RefDrone annotation process with RDAnnotator. Multiple specialized LVLM-based modules collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
@@ 2.2 Referring expression comprehension methods
REC methods can be broadly categorized into large vision language models (LVLMs) and specialist models. LVLMs [you2023ferret, chen2023minigpt, chen2023shikra, bai2023qwen, peng2023kosmos, lin2023sphinx, zhan2025griffon, wang2023cogvlm, jiang2025detect] have recently been applied to REC tasks as part of evaluating their broader visual-language understanding capabilities. These models leverage extensive referring instruction tuning data to achieve competitive performance without task-specific architectural designs. To manage computational demands, LVLMs typically process downsampled images [llava] or employ visual token reduction strategies [bai2023qwen, wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels], which limits their ability to preserve fine-grained details. This constraint makes them less effective for grounding small objects, a crucial requirement in applications like aerial image analysis.
Specialist models include two-stage and one-stage methods. Two-stage methods [hu2017modeling, liu2019learning, hong2019learning, reclip, han2024zero] typically approach REC as a ranking task: first generating region proposals through an object detector, then ranking proposals based on language-vision alignment. Despite achieving strong accuracy, two-stage pipelines suffer from slow inference. In contrast, one-stage methods [mdetr, glip, liu2023dq, gan2020large, yan2023universal, gdino] directly predict target regions guided by language input. These approaches leverage transformers to enable cross-modal interactions between visual and textual features. Among these, GroundingDINO (GDINO) [gdino] has gained widespread attention for its impressive results in REC tasks. Our work extends GDINO by incorporating explicit number modeling to handle multi-target and no-target scenarios, which are critical challenges in drone-based REC tasks.
@@ 3.1 Data source
The RefDrone benchmark is built upon VisDrone2019-DET [visdrone], a high-quality drone-captured object detection dataset. Images are collected across scenarios, illumination conditions, and flying altitudes. To ensure sufficient complexity for referring expression annotation, we filtered the dataset, retaining only images with at least three objects and excluding objects with bounding box areas smaller than 64 pixels. VisDrone2019-DET provides object categories and bounding box coordinates, which we convert to normalized center points (range 0-1). This approach reduces the token count for LVLMs while preserving spatial relationships.
@@ 3.2 RDAnnotator for semi-automated annotation
To construct our benchmark, we introduce RDAnnotator, a semi-automated annotation framework that balances annotation quality and scalability by integrating LVLM modules with a human-in-the-loop validation process. As illustrated in Figure 2, the framework operates in five steps.
Step 1: scene understanding. A scene-parsing module, powered by GPT-4o, generates three diverse captions for each image. These captions establish a foundational context by detailing spatial arrangements and object relationships for subsequent referring expression generation.
Step 2: color categorization. To identify color attributes, which are crucial discriminative features in REC, we employ a hybrid pipeline. This pipeline combines a CNN-based classifier (WideResNet-101 [wideresnet]) with high-level LVLM-based semantic reasoning to ensure accurate color attribution despite challenging lighting and atmospheric conditions.
Step 3: expression generation. We reformulate expression generation as an object grouping task. This module clusters semantically related objects and generates a textual justification for each grouping, which then serves as the referring expression. A dynamic feedback loop connects this step to color categorization (Step 2), triggering re-analysis if novel color terms are generated, thereby ensuring attribute consistency across the dataset.
Step 4: quality evaluation. A validation module automatically assesses each generated object-expression pair for semantic accuracy and uniqueness. Correct annotations advance to human verification (Step 5). Incorrect annotations are routed back with targeted feedback: semantic errors are returned to expression generation (Step 3), while color inaccuracies are sent back to color categorization (Step 2).
Figure 3: Object number distribution per expression in gRefCOCO [grefcoco] and RefDrone datasets.
Figure 4: Object size distribution analysis. (a) Object size distribution in RefDrone dataset (small: $32^{2}=1024$ pixels, normal: 1024 to 9216 pixels, large: $96^{2}=9216$ pixels). (b) Object size histograms in RefDrone and gRefCOCO [grefcoco] datasets.
Step 5: human verification. All generated annotations undergo a final human review, categorized into three tiers:
@@
  • Refinement required. Annotations with minor errors are corrected through human editing.
  • Significant issues. Annotations with poorly grounded or inconsistent content trigger full regeneration.
This multi-level feedback system ensures high annotation fidelity. Expressions that consistently fail verification are designated as “no-target” samples, creating a set of contextually valid negative examples.
Each step leverages LVLMs via in-context learning with carefully designed task-specific prompts and examples (see Appendix). The effectiveness of our framework is demonstrated by the annotation outcomes: 42% of samples were directly accepted, 47% required minor refinement, and only 11% were rejected for re-annotation. Ultimately, RDAnnotator reduces human annotation effort by 85% , decreasing the average time per expression from 7 minutes to 1 minute. This efficiency is achieved at a low API cost of \$0.0539 per expression, demonstrating the framework’s scalability for creating large-scale, high-fidelity REC datasets.
@@
Figure 6: Architecture of a single decoder layer in Number GroundingDINO. Key modifications from GDINO [gdino] (highlighted in yellow box) include: (1) a number prediction head (FFN) to estimate target count, (2) number-queries selected through the predicted number, (3) a number cross-attention between selected number-queries and detection-queries.
@@ 3.3 Dataset analysis
The RefDrone dataset comprises 17,900 referring expressions for 63,679 object instances across 8,536 images, spanning 10 categories. The dataset preserves the original train, validation, and test splits from VisDrone2019-DET [visdrone]. On average, each expression is 9.0 words long and refers to 3.8 objects. RefDrone is characterized by three primary challenges, illustrated in Figure 1:
1) Multi-target and no-target samples. In contrast to datasets like RefCOCO [refcoco] that focus on single-object references, RefDrone features a significant portion of complex queries, including 11,362 multi-target and 847 no-target expressions. The number of targets per expression ranges from 0 to 242. As shown in Figure 3, this distribution presents a greater challenge than that of gRefCOCO [grefcoco], where expressions typically refer to only one or two objects.
2) Multi-scale and small-scale target detection. The dataset exhibits a wide distribution of object scales (Figure 4): small objects ($32^{2}$ pixels) account for 31%, medium objects ($32^{2}$-$96^{2}$ pixels) for 55%, and large objects ($96^{2}$ pixels) for 14%. The high variance in object scales, particularly the prevalence of small objects, underscores multi-scale and small-scale target detection challenges.
3) Complex environment with rich contextual reasoning. Images are captured in complex environments with diverse viewpoints, lighting, and dense backgrounds. Consequently, the referring expressions extend beyond simple attributes (e.g., color) and spatial relationships (e.g., ‘left of’) to describe complex object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’). This complexity is visualized in the word clouds in Figure 5.
3.4 Comparison to existing datasets
Table 1 situates RefDrone among existing REC datasets. Key distinguishing features of RefDrone are its high average number of targets per expression and the use of an LVLM annotation pipeline. This pipeline generates expressions with richer contextual details compared to those from template-based methods or human-only annotation [zhang2023multi3drefer]. Although RIS-CQ [riscq] also uses an LLM for generation, its process is decoupled from visual content, leading to expressions that can be linguistically complex but visually ambiguous. While RSVG [zhan2023rsvg] targets small objects, the descriptive quality of its expressions is limited. In contrast, RefDrone comprehensively integrates these challenges, establishing it as a challenging benchmark in REC tasks.